# LSR-MCTS: Alleviating Long Range Dependency in Code Generation

### v3

@@ Abstract
The emergence of large language models (LLMs) has significantly promoted the development of code generation task, sparking a surge in pertinent literature. Current research is hindered by redundant generation results and a tendency to overfit local patterns in the short term. Although existing studies attempt to alleviate the issue by adopting a multi-token prediction strategy, there remains limited focus on choosing the appropriate processing length for generations. By analyzing the attention between tokens during the generation process of LLMs, it can be observed that the high spikes of the attention scores typically appear at the end of lines. This insight suggests that it is reasonable to treat each line of code as a fundamental processing unit and generate them sequentially. Inspired by this, we propose the LSR-MCTS algorithm, which leverages MCTS to determine the code line-by-line and select the optimal path. Further, we integrate a self-refine mechanism at each node to enhance diversity and generate higher-quality programs through error correction. Extensive experiments and comprehensive analyses on three public coding benchmarks demonstrate that our method outperforms the state-of-the-art performance approaches [^1] The code will be open source..
LSR-MCTS: Alleviating Long Range Dependency in Code Generation
Tingwei Lu1, Yangning Li1,2, Liyuan Wang 3, Binghuai Lin3, Jiwei Tang1, Qingsong Lv1, Wanshi Xu4, Hai-Tao Zheng1,2thanks: Corresponding author: zheng.haitao@sz.tsinghua.edu.cn, Yinghui Li1, Xin Su3, Zifei Shan3, 1Tsinghua University 2Peng Cheng Laboratory 3Tencent Technology Co., Ltd 4Peking University ltw23@mails.tsinghua.edu.cn
@@ 1 Introduction
Large language models (LLMs) such as LLaMA Touvron et al. (2023) and GPT-4 Achiam et al. (2023) have achieved tremendous success across various domains recently, particularly in NLP Team et al. (2023); Jiang et al. (2023, 2024). The code generation task aims to automatically generate code meeting the requirements based on the provided natural language (NL) description, which can be regarded as a text sequence. Thus, code generation can still be considered a specialized form of text generation, with the emergence of LLMs tailored to coding, known as Code LLMs.
Figure 1: Examples of code generated by two kinds of methods. The token-by-token approach misunderstands the NL description (highlighted in red), leading to generating a verbose program that passes only part of the test cases.
The research on Code LLMs is divided into two prime avenues: (1) Pre-train fundamental Code LLMs. Pre-trained models such as CodeT5 Wang et al. (2021), CodeGen Nijkamp et al. (2022), StarCode Li et al. (2023), and DeepSeek-Coder Guo et al. (2024) provide solid backbone for code tasks; (2) Design decoding strategy. Numerous decoding strategies Zhang et al. (2023); Zhu et al. (2024) are proposed to correct errors generated by greedy decoding during inference. These methods are promising for their plug-and-play manner. We focus on them in this paper.
Existing methods primarily generate code token-by-token using LLMs Zhang et al. (2023); Brandfonbrener et al. (2024), which pay more attention to short-term tokens at each generation. However, due to the strict logical structure and closely related knowledge inherent in programming languages, overlooking the long-term dependency on code may lead to severely flawed programs. Therefore, the token-level approaches, which concentrate on local code segments are prone to misalign code fragments with the natural language (NL) description or produce redundant among the long-term perspective. As depicted in Figure 1, the program generated token-by-token is executable but contains intricate and unnecessary thoughts that lead to errors in some test cases.
@@
To overcome the short-term issue, Gloeckle et al. (2024) explores to introduce multi-tokens prediction as auxiliary training task, which encourages LLMs to consider longer-term dependencies within the generated sequence. The paper highlights the significance of attention between distant tokens for LLMs. Inspired by it, we take a deeper dive into the attention between tokens of existing LLMs to reveal the essential connection between them. As shown in Figure 2, it is observed that certain tokens have a profound influence on subsequent generations. It can be inferred that these tokens can summarize information from the prior code and lead the following generation, which is denoted as “summary tokens” by us. Thus, ensuring the correctness of the previous summary token and the corresponding line is crucial, which is beneficial for future generations and rectification. The observation highlights that the line, rather than the token, emerges as a more effective fundamental processing unit in the code generation task.
Motivated by this, we introduce a novel decoding strategy Line-level Self-Refine Monte Carlo Tree Search, termed LSR-MCTS. It combines the line-level concept with MCTS, where each node in the tree signifies a line segment. A trajectory from the root node to the leaf forms a complete program. LSR-MCTS shortens the distance between tokens in the tree from a higher horizon and encourages the model to predict from a global optimization, generating more concise programs depicted in Figure 1 and alleviating the long-term dependency problem. Given the inherently vast search space of code generation, it is necessary to limit the number of children for each node. However, the constraint may overlook some viable branches. Considering that summary tokens can facilitate code correction, we integrate a self-refine mechanism at each node to regenerate the current line and summary token, thereby increasing the number of high-quality child nodes and ultimately enhancing performance.
Our principal contributions are as follows:
  • We propose a line-level MCTS approach that treats individual lines of code as basic processing units, enhancing code generation with a long-term perspective.
  • Integrating a self-refine mechanism into each node allows us to search a wider array of potential solutions to identify the most effective ones. Concurrently, it is conducive to rectifying lines and summary tokens.
  • Extensive experiments and comprehensive analysis on renowned coding benchmarks including HumanEval, MBPP, and Code Contests validate the exceptional performance of the LSR-MCTS decoding strategy.
@@ 2.1 LLMs for Code
LLMs have demonstrated remarkable capabilities in handling tasks such as NLP Team et al. (2023); Anil et al. (2023); Jiang et al. (2023, 2024), with their prowess particularly evident in the domain of code generation. Models like Codex Chen et al. (2021), trained across a multitude of programming languages and billions of lines of code, have emerged as versatile code snippet generators, integrated into tools like Copilot to assist programmers in coding. AlphaCode Li et al. (2022), which is trained on a vast array of open-source Python code, stands out as the first LLM capable of generating structured code directly from NL descriptions.
Stimulated by these pioneering efforts, many researchers are dedicated to the training of Code LLMs. Google introduces the proprietary PaLM-Coder Chowdhery et al. (2023), which generates code results through API calls, showcasing impressive performance. Concurrently, other researchers focus on developing open-source Code LLMs, such as Salesforce’s CodeGen Nijkamp et al. (2022), Meta’s In-Coder Fried et al. (2022), Code Llama Roziere et al. (2023), and others including StarCoder Li et al. (2023), CodeGeeX Zheng et al. (2023), DeepSeek-Coder Guo et al. (2024), etc. These models are progressively approaching and surpassing the performance of general models, bolstering the confidence in training Code LLMs and amplifying code generation efficiency.
@@ 2.2 Monte Carlo Tree Search
The performance improvement of LLMs on various tasks is attributable to not merely their augmented capabilities from training, but also the optimization of their generation strategies Yasunaga et al. (2023); Chuang et al. (2023); Huang et al. (2024). MCTS, as one of the efficient strategies for handling large-scale search spaces, is highly applicable in the generation domain and is becoming a research hotspot in code generation.
VerMCTS Brandfonbrener et al. (2024) designs a logical verifier within the MCTS process, expanding tokens until the verifier can return a score. Furthermore, PG-TD Zhang et al. (2023) proposes an MCTS-based method evaluated by test cases for code generation, treating each token decoded by LLMs as an action. However, the application of MCTS in code generation is predominantly token-level, focusing on short-term predictions, which causes a local optimal solution, especially when dealing with programs that consist of thousands of tokens. As the distance between nodes increases with the length of the code, the practicality diminishes considerably.
Figure 3: The framework of LSR-MCTS. The red part in (a) shows the four iterative steps of LSR-MCTS: selection, expansion, evaluation, and backpropagation. The green sections reflect the self-refine process, where new nodes are generated in the expansion step, and a higher-quality refined node is constructed in conjunction with LLMs and the new nodes. Part (b) more explicitly displays the content of a single node, including context, line, and supplement, with the main body "line" emphasized in bold blue. However, the first two codes are incorrect; through self-refine, they can be adjusted to the correct program.
@@ 2.3 Self-Refine Strategy
In addition to MTCS, self-refine can also be considered an efficient strategy Yao et al. (2023); Madaan et al. (2023). LATS Zhou et al. (2023) employs a generation strategy that combines MCTS with self-reflection and environmental feedback, generating multiple programs from the same node and using prompts to reflect on incorrect code predictions. Reflexion Shinn et al. (2024) continuously refines and regenerates the code based on the environment—the textual feedback from LLMs on the generated code—ceasing until the evaluation metrics reach a plateau.
@@
3 Task Formulation
Code generation is a subtask of text generation. It takes the natural language describing a problem as input, denoted as $T=\left(t_{1},t_{2},\ldots,t_{m}\right)$, and generates the code that solves the problem as output, represented as $C=\left(c_{1},c_{2},\ldots,c_{n}\right)$, where $m,n$ is the length of input and output, $t_{i},c_{j}\in\mathcal{V}$. Here, $\mathcal{V}$ is the whole vocabulary. It is a sequence-to-sequence process that relies on the model parameters $M_{\theta}$. Typically, the next token’s probability distribution is predicted based on the preceding content, and the top-k most probable tokens are sampled to iteratively generate and construct the complete code block $C$. $C$ is not only a textual sequence but also a function that can take input data in the correct format to produce output.
For the correctness of the generated code, test cases $\left[I,O\right]=\{\left(i_{j},o_{j}\right)\}_{j=1}^{J}$ are provided, where $J$ is the number of test cases. They are often divided into public and private test cases (some datasets don’t provide this division). LLMs have the option to use the former to assist generation, thereby improving the quality of the generated code, while the latter is used to assess the correctness of the code. The code is considered correct only if it satisfies $C(i_{j})=o_{j}$ for all private test cases $(i_{j},o_{j})$.
4 Method
In this section, we elaborate on the proposed training-free decoding strategy LSR-MCTS. The code is segmented into granular line-level code snippets, and each line of the code along with its pertinent information is treated as an individual node. The optimal branch node is selected through MCTS, for generating the most efficacious code block on a line-by-line basis. For each node, we introduce a self-refine mechanism to ensure a more exhaustive exploration of the solution space, leading to the emergence of higher-quality rectified code. The comprehensive depiction of the entire process is shown in Figure 3, with an elaborated pseudocode presented in Algorithm 1.
@@ 4.1 Line-level MCTS
MCTS approaches code generation task as a meticulous process of tree searching. The root node lies the initial Natural Language prompt describing the problem, with each subsequent node representing an extension of the code sequence generation. The search space is encompassed by all conceivable branches of the tree. The objective of MCTS is to navigate this potentially boundless search space, identifying the optimal child nodes to construct a coherent path that culminates in the complete code. Our LSR-MCTS framework adheres to the conventional MCTS algorithm’s four-phase structure—selection, expansion, evaluation, and backpropagation. It enriches each phase with a line-level concept, enhancing the precision of the search.
In contrast to general token-level MCTS methods, which treat each token as a node Zhang et al. (2023); DeLorenzo et al. (2024), we redefine the node information within the tree structure to suit our line-level decoding process. As illustrated in Figure 3(b), a node encompasses three components: context, line, and supplement, which together form a segment of complete and executable code. The line represents a specific line of code that characterizes the node, while the context is an n-line code block constructed from the path of ancestor nodes. To ensure that each node can be evaluated and scored using public test cases, incomplete code blocks must be supplemented to ensure completeness, hence the inclusion of the supplement component to round out the code. Here, both the line and supplement are generated by LLMs, with the next line of the current node being selected as the line, and the rest as the supplement.
During the selection phase, we apply the upper confidence bound for trees(UCT) strategy, starting from the root node to identify the most promising branch for exploration, and extending to the leaf node (Algorithm 1, Line 3-6). For node $s$, the UCT score is calculated as follows:
@@
where $s.values$ is the cumulative score of node $s$ affected by the reward scores of descendant nodes and backpropagation process, $s.visits$ is the count of times node $s$ has been visited, and $N$ represents the total number of rollouts that have been executed. The function $UCT(\cdot)$ in Algorithm 1 Line 5 means selecting the node with the highest UCT score in the children list.
In the subsequent expansion phase, LLM is utilized to generate $m$ code block $\left[C_{1},C_{2},\ldots,C_{m}\right]$ for leaf in non-terminating states, segmenting the complete code block $C_{i}$ into context, line, and supplement (Algorithm 1, Line 7-16). The line here denotes the immediate subsequent line of code following the current node’s line, and $m$ is set to 3 for constraining the number of child nodes. Concat($node$, $code$) means updating the context, line, and supplement of $node$ according to the content of $code$. $M_{\theta}(text,m)$ represents using the $text$ as input for the LLM with parameters $\theta$, and generating $m$ outputs in parallel. The content about the generation prompt of $next\_codes$ is introduced in Appendix A.1.
Upon acquiring the details of the next node, it is appended as a child to the current node. The quality of the newly generated program is then appraised using public test cases. For the code block $C$, the reward formula is defined:
$Reward(C)=\frac{b}{a}$ (2)
where $a$ denotes the total number of public test cases, and $b$ signifies the number of test cases that the code block $C$ successfully passes (Algorithm 1, Line 17-19).
Once the reward score of the program is determined, it is retroactively disseminated to the root node, updating the value and the visited count of each ancestor node along the path, accordingly promoting future decision-making (Algorithm 1, Line 20-25).
@@
0:    $M_{\theta}$: LLM with parameters $\theta$; $root$: the root node of the tree; $m$: the maximum number of children of any node; $n$: the number of max rollouts; $PR$: the general generation prompt; $SRP$: self-refine PROMPT; $R(\cdot)$: reward function for code according to the public test cases.
@@
10:        $next\_node\leftarrow$ Concat$(node,next\_code)$;
11:        Add $next\_node$ to the children of $node$;
@@
14:        $refined\_node\leftarrow$ Concat$(node,refined\_code)$;
15:        Add $refined\_node$ to the $node.children$;
@@
22:        $node.values+=r\_next+r\_refine$
@@ 4.2 Self-Refine Mechanism
To address potential omissions of feasible branches due to the limitation on the number of child nodes, and to rectify the code to guarantee the precision of summary tokens that exert substantial influence on later generations, we introduce a self-refine mechanism during the expansion phase of MCTS. This mechanism generates a new code block for each node, which serves as an unconstrained child node of the current node, allowing for further expansion in the following operations.
Due to the constraints imposed by hyperparameter settings in the line-level MCTS, code generation may be confined within a specific search space, potentially overlooking many viable search areas. Consequently, during the self-refine process, LLMs use the information within each node as a guide to explore under-searched spaces and identify high-quality child nodes, as depicted in Figure 3(a). Given that previous nodes have achieved high quality through multiple iterations of line-level MCTS, we only need to consider the subsequent generation results of the current node. Therefore, the refined node acts solely as a child of the current node, rectifying the current line and ensuring the generation of high-quality in the following steps. It utilizes a carefully crafted prompt, denoted as $SRP$, to input into LLMs for generating code, showing in Algorithm 1 Line 13.
Once the new refined nodes are obtained, they are processed in the same manner as regular nodes during the evaluation and backpropagation stages(Algorithm 1 Line 17-25), undergoing quality assessment and backpropagation in sequence to ultimately generate a superior program.
@@ 5 Experiments
In this section, extensive experiments are conducted to substantiate the effectiveness of our method and its superiority over existing technologies. The experimental setup includes the selection of datasets, models, baselines, and evaluation metrics, as well as a detailed exposition of the experimental procedures and results.
@@
Table 1: Main results of code generation performance on three public benchmarks. LSR-MCTS is compared with other decoding strategies such as Beam-Search, Reflexion, and PG-TD. The best results are highlighted in light blue.

### v4

# From Token to Line: Enhancing Code Generation with a Long-Term Perspective

@@ Abstract
The emergence of large language models (LLMs) has significantly promoted the development of code generation task, sparking a surge in pertinent literature. Current research is hindered by redundant generation results and a tendency to overfit local patterns in the short term. Although existing studies attempt to alleviate the issue by adopting a multi-token prediction strategy, there remains limited focus on choosing the appropriate processing length for generations. By analyzing the attention between tokens during the generation process of LLMs, it can be observed that the high spikes of the attention scores typically appear at the end of lines. This insight suggests that it is reasonable to treat each line of code as a fundamental processing unit and generate them sequentially. Inspired by this, we propose the LSR-MCTS algorithm, which leverages MCTS to determine the code line-by-line and select the optimal path. Further, we integrate a self-refine mechanism at each node to enhance diversity and generate higher-quality programs through error correction. Extensive experiments and comprehensive analyses on three public coding benchmarks demonstrate that our method outperforms the state-of-the-art performance approaches.
Index Terms—  Code Generation, Monte Carlo Tree Search, Line-level Decoding
@@ 1 Introduction
Large language models (LLMs) such as LLaMA [31] and GPT-4 [1] have achieved tremendous success across various domains recently, particularly in NLP [30, 23, 18, 6, 17]. The code generation task aims to automatically generate code meeting the requirements based on the provided natural language (NL) description, which can be regarded as a text sequence. Thus, code generation can still be considered a specialized form of text generation, with the emergence of LLMs tailored to coding, known as Code LLMs.
Fig. 1: Examples of code generated by two kinds of methods.
@@
The research on Code LLMs is divided into two prime avenues: (1) Pre-train Code LLMs. Pre-trained models such as CodeGen [22], StarCode [16], and DeepSeek-Coder [12] provide solid backbone for code tasks; (2) Design decoding strategy. Numerous decoding strategies [34, 39] are proposed to correct errors generated by greedy decoding during inference. These methods are promising for their plug-and-play manner. We focus on the second in this paper.
Existing methods primarily generate code token-by-token using LLMs [34, 4], which pay more attention to short-term tokens at each generation. However, due to the strict logical structure and closely related knowledge inherent in programming languages, overlooking the long-term dependency on code may lead to severely flawed programs. Therefore, the token-level approaches, which concentrate on local code segments, are prone to misaligning code fragments with the natural language (NL) description or producing redundancy among the long-term perspective, as depicted in Figure 1.
To overcome the short-term issue, Gloeckle et al. [11] explore introducing multi-tokens prediction as an auxiliary training task, which encourages LLMs to consider longer-term dependencies within the generated sequence. The paper highlights the significance of attention between distant tokens for LLMs. Following this work, we take a deeper dive into the attention between tokens of existing LLMs. As shown in Figure 2, it is observed that certain tokens have a profound influence on subsequent generations. It can be inferred that these tokens can summarize information from the prior code and lead the following generation, which is denoted as “summary tokens”. Thus, ensuring the correctness of the previous summary token and the corresponding line is crucial, which is beneficial for future generations and rectification. The observation highlights that the line emerges as a more effective fundamental processing unit in the code generation task.
Motivated by this, we introduce a novel decoding strategy Line-level Self-Refine Monte Carlo Tree Search, termed LSR-MCTS. It combines the line-level concept with MCTS, where each node in the tree signifies a line segment. A trajectory from the root to the leaf forms a complete program. LSR-MCTS shortens the distance between tokens in the tree from a higher horizon and encourages the model to predict from a global optimization, generating more concise programs depicted in Figure 1. However, the higher horizon may overlook some viable branches. Considering that summary tokens can facilitate code correction, we integrate a self-refine mechanism at each node to regenerate the current line and summary token. Extensive experiments and comprehensive analysis validate the exceptional performance of the LSR-MCTS decoding strategy.
@@ 2.1 LLMs for Code
LLMs have demonstrated remarkable capabilities in handling tasks such as NLP [29, 28, 2, 14, 15]. Numerous studies have shown that excessively long input text leads to performance degradation [26, 27, 20, 35, 36]. As for the field of code generation, models like Codex [5], trained across a multitude of programming languages and billions of lines of code, have emerged as versatile code snippet generators, integrated into tools like Copilot to assist programmers in coding. AlphaCode [19], which is trained on a vast array of open-source Python code, stands out as the first LLM capable of generating structured code directly from NL descriptions.
Stimulated by these pioneering efforts, many researchers are dedicated to the training of Code LLMs. Google introduces the proprietary PaLM-Coder [7], which generates code results through API calls, showcasing impressive performance. Concurrently, other researchers focus on developing open-source Code LLMs, such as Salesforce’s CodeGen [22], Meta’s In-Coder [10], Code Llama [24], and others including StarCoder [16], CodeGeeX [37], DeepSeek-Coder [12], etc. These models are progressively approaching and surpassing the performance of general models, bolstering the confidence in training Code LLMs and amplifying code generation efficiency.
Fig. 3: The framework of LSR-MCTS. The red part in (a) shows the four iterative steps of LSR-MCTS: selection, expansion, evaluation, and backpropagation. The green sections reflect the self-refine process, where new nodes are generated in the expansion step. Part (b) explicitly displays the content of a single node, including context, line, and supplement, with the main body ”line” emphasized in bold blue.
@@ 2.2 Monte Carlo Tree Search
The performance improvement of LLMs on various tasks is attributable to not merely their augmented capabilities from training, but also the optimization of their generation strategies [33, 8, 13]. MCTS, as one of the efficient strategies for handling large-scale search spaces, is highly applicable in the generation domain and is becoming a research hotspot in code generation.
VerMCTS [4] designs a logical verifier within the MCTS process, expanding tokens until the verifier can return a score. Furthermore, PG-TD [34] proposes an MCTS-based method evaluated by test cases for code generation, treating each token decoded by LLMs as an action. However, the application of MCTS in code generation is predominantly token-level, focusing on short-term predictions, which causes a local optimal solution, especially when dealing with programs that consist of thousands of tokens. As the distance between nodes increases with the length of the code, the practicality diminishes considerably.
@@ 2.3 Self-Refine Strategy
In addition to MTCS, self-refine can also be considered an efficient strategy [32, 21]. LATS [38] employs a generation strategy that combines MCTS with self-reflection and environmental feedback, generating multiple programs from the same node and using prompts to reflect on incorrect code predictions. Reflexion [25] continuously refines and regenerates the code based on the environment—the textual feedback from LLMs on the generated code—ceasing until the evaluation metrics reach a plateau.
@@
3 Methodology
In this section, we elaborate on the proposed training-free decoding strategy LSR-MCTS. The comprehensive depiction of the entire process is shown in Figure 3.
@@ 3.1 Line-level MCTS
MCTS treats code generation task as a meticulous process of tree search. The root node lies the initial Natural Language prompt describing the problem, with each subsequent node representing an extension of the generated code. The search space encompasses all conceivable branches of the tree. The objective of MCTS is to navigate this potentially unbounded search space by identifying the optimal child nodes, gradually constructing a coherent path that culminates in the complete and correct code solution.
LSR-MCTS framework adheres to the conventional MCTS algorithm’s four-phase structure—selection, expansion, evaluation, and backpropagation. It enriches each phase with a line-level concept, enhancing the precision of the search. We redefine the node information within the tree structure to suit our line-level decoding process. As illustrated in Figure 3(b), a node encompasses context, line, and supplement, which together form a segment of complete and executable code. The line represents a specific line of code that characterizes the node, while the context is an n-line fixed code block constructed from the path of ancestor nodes. To ensure that each node can be evaluated using public test cases, incomplete code blocks must be supplemented. Here, both the line and supplement are generated by LLMs, with the next line of the current node being selected as the line, and the rest as the supplement.
During the selection phase, we apply the upper confidence bound for trees(UCT) strategy, starting from the root node to the max score leaf node $s$:
@@
where $s.values$ is the cumulative score of node $s$, $s.visits$ is the count of times node $s$ has been visited, and $N$ represents the number of rollouts that have been executed. In the subsequent expansion phase, LLM is utilized to generate $m$ code block for leaf in non-terminating states, segmenting the complete code block $C$ into context, line, and supplement. Here $m$ is set to 3 for constraining the number of child nodes. Upon acquiring the details of the next node, it is appended as a child to the current node. The quality of the generated code is then evaluated by the pass rate of public test cases. Once the score of the code is determined, it is backpropagated to the root node, updating the value and the visited count of each ancestor node along the path, accordingly promoting future decision-making. To improve the efficiency of code generation, we add a cache mechanism to store the code blocks of ancestor nodes.
@@
[…]
@@
10:    $next\_node\leftarrow$ Concat(node,next_code)(node,next\_code);
11:    Add next_nodenext\_node to the children of nodenode;
@@
14:    $refined\_node\leftarrow$ Concat(node,refined_code)(node,refined\_code);
15:    Add refined_noderefined\_node to the $node.children$;
@@
22:    node.values+=r_next+r_refinenode.values+=r\_next+r\_refine
@@ 3.2 Self-Refine Mechanism
To address omissions of feasible branches due to the limitation on the number of child nodes, and to rectify the code to guarantee the precision of summary tokens that exert substantial influence on later generations, we introduce a self-refine mechanism during the expansion phase of Line-level MCTS.
This mechanism generates a new code block for any node whose score is low and for the last node in the current path. The newly created block acts as an unconstrained child of the current node, enabling further expansion in subsequent operations. The score threshold is empirically set to 0.5 to balance exploration and efficiency. Once the new refined nodes are obtained, they are processed in the same manner as regular nodes during the evaluation and backpropagation stages, ultimately generating a superior program. It can be seen in Figure 3 (b) that the first two codes are incorrect. Through self-refine, they can be adjusted to the correct code.
@@ 4 Experiments
Table 1: Main results and ablation performance on three public benchmarks. The best results are highlighted in light blue.